The sum of the digits of a two-digit counting number is 10. If the digits are reversed, the new number is one less that twice the original number. What is the original number?
(please help by showing work.)

Answer by oscargut(2103)

Let the number "ab"
Reversed number = "ba"
a+b=10
10b+a = 2(10a+b)-1
10b+a = 20a+2b-1
10b+a -20a -2b +1 =0
-19a+8b+1=0
Since b = 10-a
-19a+8(10-a)+1=0
-19a+80-8a+1=0
-27a=-81
a = 3
b=7
Answer: original number is 37
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com